# Q3R: Quadratic Reweighted Rank Regularizer for Effective Low-Rank Training

### v1

@@ Abstract
Parameter-efficient training, based on low-rank optimization, has become a highly successful tool for fine-tuning large deep-learning models. However, these methods fail at low-rank pre-training tasks where maintaining the low-rank structure and the objective remains a challenging task. We propose the Quadratic Reweighted Rank Regularizer dubbed Q3R, which leads to a novel low-rank inducing training strategy inspired by the iteratively reweighted least squares (IRLS) framework. Q3R is based on a quadratic regularizer term which majorizes a smoothed log determinant serving as rank surrogate objective. Unlike other low-rank training techniques, Q3R is able to train weight matrices with prescribed, low target ranks of models that achieve comparable predictive performance as dense models, with small computational overhead, while remaining fully compatible with existing architectures. For example, we demonstrated one experiment where we are able to truncate $60\%$ and $80\%$ of the parameters of a ViT-Tiny model with $1.3\%$ and $4\%$ accuracy drop in CIFAR-10 performance respectively. The efficacy of Q3R is confirmed on Transformers across both image and language tasks, including for low-rank fine-tuning.
@@ 2 Contribution
In this paper, we propose the Quadratic Reweighted Rank Regularization (Q3R), which bridges this gap by introducing an optimizer-compatible regularization framework based on smoothed log-determinant rank surrogates outlined in Section˜4.1, which is specifically designed for low-rank pre-training. While theoretically grounded in saddle-escaping second-order optimization methods, it comes with little computational overhead compared to unregularized training despite its efficacy for promoting low-rank neural network weight matrices. Additionally, we propose the Adam variant AdamQ3R in Section˜4.2, which is tailored to optimizing Q3R-regularized loss functions and which improves the performance of training Q3R-regularized models.
We demonstrate in numerical experiments that Q3R is able to reduce the number of parameters in ViT models by, for example, $60\%$ during pre-training on CIFAR-10, with only around $1.3\%$ accuracy drop. We validate the performance of Q3R for low-rank fine-tuning with experiments fine-tuning RoBERTa and Llama3 on GLUE tasks, for which Q3R achieves comparable performance compared to dense fine-tuning and state-of-the art low-rank PEFT methods. Compared to state-of-the-art low-rank training methods such as LoRA [29], LoRITa [3], Q3R consistently produces models with better generalization at high truncation levels, without requiring overparameterization or full-rank warmup phases.
@@ 3 Related Work
Parameter-efficient fine-tuning is the concept of modifying only parts of a fully parametrized pre-trained model to excel at specific task of interest. PEFT methods such as adapters [25] and LoRA [29] introduce small, trainable modules into a frozen pretrained model, drastically reducing the number of parameters to be updated. These techniques often match full fine-tuning performance with only a tiny fraction of trainable parameters. However, the low-rank constraints that make LoRA-style methods efficient for downstream tasks also limit capacity if applied during pre-training. Training from scratch with only low-rank adapters or factorizations (instead of full-rank weight updates) tends to underperform, as it restricts optimization to a low-dimensional subspace [75]. LoRA assumes a well-formed pretrained weight $W$ plus a low-rank perturbation of rank-$r$ adapter matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$ such that $\Delta W=AB$; without a strong initial $W$, such updates struggle to capture the full complexity needed for learning from scratch. KronA replaces LoRA’s product with a Kronecker factorization for better rank-parameter trade-offs [18]. DoRA decouples update magnitude and direction via a learnable scaling factor, improving upon LoRA’s expressivity [45]. Compacter uses shared, low-rank, Kronecker-parameterized adapters across layers, matching standard adapters with only 0.05 % extra parameters [48].
@@
In a line of work that significantly precedes the interest in low-rank techniques for deep learning, the problem of identifying or learning a low-rank matrix from noisy, under-determined linear measurements has been studied for decades in control theory [19, 16], recommender systems [35, 31] and compressed sensing [53, 17]. Even in this setting, which is fundamentally linear unlike the training of deep neural networks, the minimization of a rank objective subject to the constraints is NP-hard [51, 53] , motivating surrogate formulations or relaxations. Convex relaxations using the nuclear norm [11, 53, 17, 12] have been popular for a long time due to their strong recovery guarantees under suitable assumptions and their ability to be tackled using the machinery of convex optimization [7], but fail to result in a convex formulation in the deep learning setting. Even disregarding computational limitations of convex regularizations [8], non-convex rank surrogates such as the log-determinant penalty [19] lead to algorithms which are more data-efficient as a evidenced for a variety of low-rank matrix recovery problems [20, 9, 37, 33].
@@ 4.1 Low-Rank Regularization via Q3R
where $\sigma_{i}(\mathbf{W})$ is the $i$-th singular value of $\mathbf{W}$, and which is parametrized by a smoothing parameter $\epsilon>0$. The definition of Equation˜1 follows [33] and is related to the log-determinant heuristics $\log\operatorname{det}(\mathbf{W}+\epsilon I)=\sum_{i=1}^{d}\log(\sigma_{i}(\mathbf{W})+\epsilon)$ defined for positive semi-definite matrices $\mathbf{W}\in\operatorname{\mathbb{R}}^{d\times d}$ of [19, 9, 52]. Compared to other log-determinant type functions, $F_{\epsilon}(\cdot)$ from Equation˜1 has a few advantages: It is lower bounded by 0 for any $\epsilon$ and has a $1$-Lipschitz gradient, making the objective compatible with gradient-based optimizers without extensive step-size adaptation. Furthermore, its smoothing parameter $\epsilon$ regulates the non-convexity of its optimization landscape and recovers the well-known squared Frobenius norm as $F_{\epsilon}(\mathbf{W})=\frac{1}{2}\|\mathbf{W}\|_{F}^{2}$ in the case of $\sigma_{1}(\mathbf{W})<\epsilon$.
@@
For $\epsilon>0$ and $\mathbf{W}^{\prime}$, let $\mathbf{U}\in\operatorname{\mathbb{R}}^{d_{1}\times r(\mathbf{W}^{\prime},\epsilon)}$, $\mathbf{V}\in\operatorname{\mathbb{R}}^{d_{1}\times r(\mathbf{W}^{\prime},\epsilon)}$ and $\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ be as in Definition˜4.1. Then the following statements are true:
@@
1:Input: NN weight matrix $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$; target rank $r_{\text{target}}$}; prev. smoothing parameter $\epsilon_{\text{old}}$.
2:Output: Updated $\epsilon_{\text{new}}$, reweighting operator $\mathcal{R}_{\mathbf{W},\epsilon_{\text{new}}}$ (via $\Sigma$, $\mathbf{U}$, $\mathbf{V}$), envelope rank $r_{\text{env}}$
3: Compute $\mathbf{U},\Sigma,\mathbf{V}]=\operatorname{SVD}^{\epsilon_{\text{old}}}(\mathbf{W})$ of $\mathbf{W}$, where $\operatorname{SVD}^{\epsilon_{\text{old}}}(\cdot)$ computes a partial singular value decomposition of its input up to order $r(\cdot,\epsilon_{\text{old}})$ (see Equation˜3) as well as $\sigma_{r_{\text{target}}+1}(\mathbf{W})$.
@@
6: Set $\mathbf{U}=\mathbf{U}_{:,1:r_{\text{env}}}$, $\mathbf{V}=\mathbf{V}_{:,1:r_{\text{env}}}$, $\Sigma=\Sigma_{1:r_{\text{env}},1:r_{\text{env}}}$ $\triangleright$ Restrict part. SVD matrices
7:return Reweighting operator $\mathcal{R}_{\mathbf{W},\epsilon_{\text{new}}}$ implicitly defined by $\mathbf{U}\!\in\!\operatorname{\mathbb{R}}^{d_{1}\times r_{\text{env}}}$, $\mathbf{V}\!\in\!\operatorname{\mathbb{R}}^{d_{2}\times r_{\text{env}}}$, $\Sigma$ & $\epsilon_{\text{new}}$.
@@
$\epsilon\leftarrow\min(\epsilon,\sigma_{r_{\text{target}}+1}(\mathbf{W}))$ (7)
@@ 4.2 Neural Network Training via AdamQ3R
Let now $y:\operatorname{\mathbb{R}}^{d_{1}^{1}\times d_{2}^{1}}\times\operatorname{\mathbb{R}}^{d_{1}^{K}\times d_{2}^{K}}\times\operatorname{\mathbb{R}}^{d_{\text{in}}}\to\operatorname{\mathbb{R}}^{d_{\text{out}}}$ be the input-output mapping of a deep neural network that depends on weight parameter matrices $\Theta=\{\mathbf{W}_{k}\in\operatorname{\mathbb{R}}^{d_{1}^{k}\times d_{2}^{k}}:\mathbf{W}_{k}\text{ is weight matrix},k=1,\ldots,K\}$. For a Transformer-based architecture such as Vision Transformer [14], the weight matrices include square layer- and head-wise query, key and value weight matrices $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\operatorname{\mathbb{R}}^{d\times d}$ as well as rectangular projection and MLP layer weight matrices. Given a pairwise loss $\ell(\cdot,\cdot)$ such as cross entropy and a training dataset $\{x_{i},y_{i}\}_{i=1}^{n}$, we can define the (unregularized) network loss as $\mathcal{L}(\Theta)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(y(\Theta,x_{i}),y_{i}\bigr)$.
@@
In order to gradually impose low-rank weights during training, we propose to optimize instead the Q3R-regularized total loss $\mathcal{L}_{\texttt{Q3R}}(\Theta):=\mathcal{L}(\Theta)+\lambda\sum_{k=1:\text{$\texttt{Q3R}$ is active for }\mathbf{W}_{k}}^{K}\operatorname{Q3R}_{\mathbf{W}_{k}^{{}^{\prime}},\epsilon_{k}}(\mathbf{W}_{k})$,
@@ 5.1 Low-Rank Pre-Training
ViT-Base Trained on CIFAR-100.
To demonstrate the performance of the low-rank pre-training methods on a more challenging dataset and larger model, we train the $86M$ parameter ViT-Base from scratch for 100 epochs on CIFAR-100 with data augmentation and $\alpha=0.0001$ [56]. In line with practice for large-scale Transformers [40], we apply low-rank training techniques solely to the multi-head self attention blocks (QKV, but not to the MLP blocks). Despite the additional size and complexity of ViT-Base compared to ViT-Tiny, Q3R remains robust and we see in Figures˜2(a) and 2(b)that it exhibits larger performance advantages with 0.400.40-$0.44$ test accuracy when $20\%$ parameters retained, whereas LoRITa models do not exceed an accuracy of $0.25$ at the same truncation level despite their substantial overparametrization. Figures˜2(a) and 2(b) are in reference to the Table˜6 in the supplementary material.
@@
We train ViT-Tiny for 100 epochs on CIFAR-10 [36] from scratch with learning rate $\alpha=0.0004$, with low-rank regularization applied only to attention weights. We evaluate the methods for a larger set of hyperparameters as shown in Figure˜3(b) using layer-wise truncation levels with retained parameter percentages $p\in\{5\%,10\%,15\%,20\%,30\%,40\%,50\%,60\%,70\%,80\%,90\%\}$, and present results in Figure˜3. Figure˜3(a) shows that Q3R experiences almost no performance drop up to $p=30\%$ for most parameter choices, exceeding the performance of reference methods. Figure˜3(b) shows that the worst performing Q3R model still outperforms any LoRA, LoRITa, or vanilla ViT-Tiny below $p=60\%$, showcasing the method’s robustness. Figures˜3 and 3(a) is reference from Table˜5 in the supplementary material.
@@ 5.2 Low-Rank Fine-Tuning
Q3R not only induces a low-rank structure during pre-training in a memory-efficient manner, but also extends naturally to compact fine‐tuning. We fine‐tune pre‐trained RoBERTa models on the GLUE benchmark using AdamQ3R with the proposed Q3R regularizer, and compare against full fine‐tuning and LoRA [29]. We impose Q3R on the weight matrices that are added to the full-rank pretrained weight matrices. For LoRA, we adopt the hyperparameters from [29], and for Q3R we cross-validate the learning rate and regularization hyperparameter $\lambda$. As shown in Table˜3, Q3R matches or exceeds LoRA’s accuracy on most tasks and exhibits a performance closer to dense fine-tuning. These results demonstrate that Q3R can serve as a unified, low‐rank training strategy—both for pre‐training and fine‐tuning of Transformer models. We discuss additional fine-tuning experiments in Section˜D.2.

### v2

@@ Abstract
Parameter-efficient training based on low-rank optimization has become a highly successful tool for fine-tuning large deep learning models. However, these methods often fail for low-rank pre-training, where simultaneously maintaining low-rank weight structure and optimizing the task objective remains challenging. We propose the Quadratic Reweighted Rank Regularizer (Q3R), which leads to a novel low-rank-inducing training strategy inspired by the iteratively reweighted least squares (IRLS) framework. Q3Ris based on a quadratic regularizer term that majorizes a smoothed log-determinant rank surrogate. Unlike other low-rank training techniques, Q3Rcan train weight matrices to prescribed low target ranks while achieving predictive performance comparable to dense models, with small computational overhead and full compatibility with existing architectures. For example, we demonstrate a Q3R-regularized ViT-Tiny experiment where truncating the model to $60\%$ and $80\%$ of its parameters results in only minor absolute accuracy drops of $1.3\%$ and $4\%$, respectively, on CIFAR-10. We confirm the efficacy of Q3Ron Transformers across both vision and language tasks, including low-rank fine-tuning.
@@ 2 Contribution
In this paper, we propose the Quadratic Reweighted Rank Regularizer (Q3R), which bridges this gap by introducing an optimizer-compatible regularization framework based on smoothed log-determinant rank surrogates outlined in Section˜4.1 and specifically designed for low-rank pre-training. While theoretically grounded in saddle-escaping second-order optimization methods, it comes with little computational overhead compared to unregularized training despite its efficacy for promoting low-rank neural network weight matrices. Additionally, we propose the Adam variant AdamQ3R in Section˜4.2, which is tailored to optimizing Q3R-regularized loss functions and which improves the performance of training Q3R-regularized models.
We demonstrate in numerical experiments that Q3R is able to reduce the number of parameters in ViT models by, for example, $60\%$ during pre-training on CIFAR-10, with only around $1.3\%$ accuracy drop. We validate the performance of Q3R for low-rank fine-tuning with experiments fine-tuning RoBERTa and Llama3 on GLUE tasks, for which Q3R achieves comparable performance compared to dense fine-tuning and state-of-the-art low-rank PEFT methods. Compared to state-of-the-art low-rank training methods such as LoRA [27], LoRITa [1], Q3R consistently produces models with better generalization at high truncation levels, without requiring overparameterization or full-rank warmup phases.
@@ 3 Related Work
Parameter-efficient fine-tuning is the concept of modifying only parts of a fully parametrized pretrained model to excel at a specific task of interest. PEFT methods such as adapters [26] and LoRA [27] introduce small, trainable modules into a frozen pretrained model, drastically reducing the number of parameters to be updated. These techniques often match full fine-tuning performance with only a tiny fraction of trainable parameters. However, the low-rank constraints that make LoRA-style methods efficient for downstream tasks also limit capacity if applied during pre-training. Training from scratch with only low-rank adapters or factorizations (instead of full-rank weight updates) tends to underperform, as it restricts optimization to a low-dimensional subspace [72]. LoRA assumes a well-formed pretrained weight $W$ plus a low-rank perturbation of rank-$r$ adapter matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$ such that $\Delta W=AB$; without a strong initial $W$, such updates struggle to capture the full complexity needed for learning from scratch. KronA replaces LoRA’s product with a Kronecker factorization for better rank-parameter trade-offs [17]. DoRA decouples update magnitude and direction via a learnable scaling factor, improving upon LoRA’s expressivity [42]. Compacter uses shared, low-rank, Kronecker-parameterized adapters across layers, matching standard adapters with only 0.05 % extra parameters [46].
@@
In a line of work that significantly precedes the interest in low-rank techniques for deep learning, the problem of identifying or learning a low-rank matrix from noisy, under-determined linear measurements has been studied for decades in control theory [18, 15], recommender systems [33, 32] and compressed sensing [53, 13]. Even in this setting, which is fundamentally linear unlike the training of deep neural networks, the minimization of a rank objective subject to the constraints is NP-hard [50, 53], motivating surrogate formulations or relaxations. Convex relaxations using the nuclear norm [9, 53, 13, 7] have been popular for a long time due to their strong recovery guarantees under suitable assumptions and their ability to be tackled using the machinery of convex optimization [5], but fail to result in a convex formulation in the deep learning setting. Even disregarding computational limitations of convex regularizations [10], non-convex rank surrogates such as the log-determinant penalty [18] lead to algorithms which are more data-efficient, as evidenced for a variety of low-rank matrix recovery problems [19, 8, 37, 35].
@@ 4.1 Low-Rank Regularization via Q3R
where $\sigma_{i}(\mathbf{W})$ is the $i$-th singular value of $\mathbf{W}$, and which is parameterized by a smoothing parameter $\epsilon>0$. The definition of Equation˜1 follows [35] and is related to the log-determinant heuristics $\log\operatorname{det}(\mathbf{W}+\epsilon I)=\sum_{i=1}^{d}\log(\sigma_{i}(\mathbf{W})+\epsilon)$ defined for positive semi-definite matrices $\mathbf{W}\in\operatorname{\mathbb{R}}^{d\times d}$ of [18, 8, 52]. Compared to other log-determinant type functions, $F_{\epsilon}(\cdot)$ from Equation˜1 has a few advantages: It is lower bounded by 0 for any $\epsilon$ and has a $1$-Lipschitz gradient, making the objective compatible with gradient-based optimizers without extensive step-size adaptation. Furthermore, its smoothing parameter $\epsilon$ regulates the non-convexity of its optimization landscape and recovers the well-known squared Frobenius norm as $F_{\epsilon}(\mathbf{W})=\frac{1}{2}\|\mathbf{W}\|_{F}^{2}$ in the case of $\sigma_{1}(\mathbf{W})<\epsilon$.
@@
For $\epsilon>0$ and $\mathbf{W}^{\prime}$, let $\mathbf{U}\in\operatorname{\mathbb{R}}^{d_{1}\times r(\mathbf{W}^{\prime},\epsilon)}$, $\mathbf{V}\in\operatorname{\mathbb{R}}^{d_{2}\times r(\mathbf{W}^{\prime},\epsilon)}$ and $\mathcal{R}_{\mathbf{W}^{\prime},\epsilon}:\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}\to\operatorname{\mathbb{R}}^{d_{1}\times d_{2}}$ be as in Definition˜4.1. Then the following statements are true:
@@
0: NN weight matrix $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$; target rank $r_{\text{target}}$}; prev. smoothing parameter $\epsilon_{\text{old}}$.
0: Updated $\epsilon_{\text{new}}$, reweighting operator $\mathcal{R}_{\mathbf{W},\epsilon_{\text{new}}}$ (via $\Sigma$, $\mathbf{U}$, $\mathbf{V}$), envelope rank $r_{\text{env}}$
1: Compute $\mathbf{U},\Sigma,\mathbf{V}]=\operatorname{SVD}^{\epsilon_{\text{old}}}(\mathbf{W})$ of $\mathbf{W}$, where $\operatorname{SVD}^{\epsilon_{\text{old}}}(\cdot)$ computes a partial singular value decomposition of its input up to order $r(\cdot,\epsilon_{\text{old}})$ (see Equation˜3) as well as $\sigma_{r_{\text{target}}+1}(\mathbf{W})$.
@@
4: Set $\mathbf{U}=\mathbf{U}_{:,1:r_{\text{env}}}$, $\mathbf{V}=\mathbf{V}_{:,1:r_{\text{env}}}$, $\Sigma=\Sigma_{1:r_{\text{env}},1:r_{\text{env}}}$ $\triangleright$ Restrict part. SVD matrices
5:return Reweighting operator $\mathcal{R}_{\mathbf{W},\epsilon_{\text{new}}}$ implicitly defined by $\mathbf{U}\!\in\!\operatorname{\mathbb{R}}^{d_{1}\times r_{\text{env}}}$, $\mathbf{V}\!\in\!\operatorname{\mathbb{R}}^{d_{2}\times r_{\text{env}}}$, $\Sigma$ & $\epsilon_{\text{new}}$.
@@
$\epsilon\leftarrow\min\bigl(\epsilon,\sigma_{r_{\text{target}}+1}(\mathbf{W})\bigr)$ (7)
@@ 4.2 Neural Network Training via AdamQ3R
Let now $y:\Pi_{\ell=1}^{K}\operatorname{\mathbb{R}}^{d_{1}^{\ell}\times d_{2}^{\ell}}\times\operatorname{\mathbb{R}}^{d_{\text{in}}}\to\operatorname{\mathbb{R}}^{d_{\text{out}}}$ be the input-output mapping of a deep neural network that depends on weight parameter matrices $\Theta=\{\mathbf{W}_{k}\in\operatorname{\mathbb{R}}^{d_{1}^{k}\times d_{2}^{k}}:\mathbf{W}_{k}\text{ is weight matrix},k=1,\ldots,K\}$. For a Transformer-based architecture such as Vision Transformer [16], the weight matrices include square layer- and head-wise query, key and value weight matrices $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\operatorname{\mathbb{R}}^{d\times d}$ as well as rectangular projection and MLP layer weight matrices. Given a pairwise loss $\ell(\cdot,\cdot)$ such as cross entropy and a training dataset $\{x_{i},y_{i}\}_{i=1}^{n}$, we can define the (unregularized) network loss as $\mathcal{L}(\Theta)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(y(\Theta,x_{i}),y_{i}\bigr)$.
@@
In order to gradually impose low-rank weights during training, we propose to optimize instead the Q3R-regularized total loss
$\mathcal{L}_{\texttt{Q3R}}(\Theta)\coloneqq\mathcal{L}(\Theta)+\lambda\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}k=1\\ \text{$\texttt{Q3R}$ is active for }\mathbf{W}_{k}\end{subarray}}^{K}\!\!\!\!\!\!\!\!\!\operatorname{Q3R}_{\mathbf{W}_{k}^{{}^{\prime}},\epsilon_{k}}(\mathbf{W}_{k}),$ (8)
@@ 5.1 Low-Rank Pre-Training
ViT-Base Trained on CIFAR-100.
To demonstrate the performance of the low-rank pre-training methods on a more challenging dataset and larger model, we train the $86M$ parameter ViT-Base from scratch for 100 epochs on CIFAR-100 with data augmentation and $\alpha=0.0001$ [56]. In line with practice for large-scale Transformers [41], we apply low-rank training techniques solely to the multi-head self-attention blocks (QKV, but not to the MLP blocks). Despite the additional size and complexity of ViT-Base compared to ViT-Tiny, Q3R remains robust and we see in Figures˜2(a) and 2(b) that it exhibits larger performance advantages with 0.400.40-$0.44$ test accuracy when $20\%$ parameters retained, whereas LoRITa models do not exceed an accuracy of $0.25$ at the same truncation level despite their substantial overparameterization. Figures˜2(a) and 2(b) are in reference to the Table˜6 in the supplementary material.
@@
We train ViT-Tiny for 100 epochs on CIFAR-10 [34] from scratch with learning rate $\alpha=0.0004$, with low-rank regularization applied only to attention weights. We evaluate the methods for a larger set of hyperparameters as shown in Figure˜3(b) using layer-wise truncation levels with retained parameter percentages $p\in\{5\%,10\%,15\%,20\%,30\%,40\%,50\%,60\%,70\%,80\%,90\%\}$, and present results in Figure˜3. Figure˜3(a) shows that Q3R experiences almost no performance drop up to $p=30\%$ for most parameter choices, exceeding the performance of reference methods. Figure˜3(b) shows that the worst performing Q3R model still outperforms any LoRA, LoRITa, or vanilla ViT-Tiny below $p=60\%$, showcasing the method’s robustness. Figures˜3 and 3(a) are based on Table˜5 in the supplementary material.
@@ 5.2 Low-Rank Fine-Tuning
Q3R not only induces a low-rank structure during pre-training in a memory-efficient manner, but also extends naturally to compact fine-tuning. We fine-tune pretrained RoBERTa models on the GLUE benchmark using AdamQ3R with the proposed Q3R regularizer, and compare against full fine-tuning and LoRA [27]. We impose Q3R on the weight matrices that are added to the full-rank pretrained weight matrices. For LoRA, we adopt the hyperparameters from [27], and for Q3R we cross-validate the learning rate and regularization hyperparameter $\lambda$. As shown in Table˜3, Q3R matches or exceeds LoRA’s accuracy on most tasks and exhibits a performance closer to dense fine-tuning. These results demonstrate that Q3R can serve as a unified, low-rank training strategy—both for pre-training and fine-tuning of Transformer models. We discuss additional fine-tuning experiments in Section˜D.2.